# Khushbu Pahwa3, Jian Ding1, Mohamed Elhoseiny1

### v3

@@ Abstract
Understanding long-form videos, such as movies and TV episodes ranging from tens of minutes to two hours, remains a significant challenge for multi-modal models. Existing benchmarks often fail to test the full range of cognitive skills needed to process these temporally rich and narratively complex inputs. Therefore, we introduce InfiniBench, a comprehensive benchmark designed to evaluate the capabilities of models in long video understanding rigorously. InfiniBench offers: (1) Over 1,000 hours of video content, with an average video length of 53 minutes. (2) The largest set of question-answer pairs for long video comprehension, totaling around 91K. (3) Eight diverse skills that span both grounding-based (e.g., scene transitions, character actions) and reasoning-based (e.g., deep context understanding, multi-event linking). (4) Rich annotation formats, including both multiple-choice and open-ended questions. We conducted an in-depth evaluation across both commercial (GPT-4o, Gemini 2.0 Flash) and most recent open-source vision-language models such as Qwen2.5-VL, InternVL3.0). Results reveal that: (1) Models struggle across the board: Even the best model, GPT-4o, achieves only 47.1% on grounding-based skills, with most models performing near or just above random chance. (2) Strong reliance on world knowledge: Models achieve surprisingly high scores using only metadata (e.g., video titles), highlighting a tendency to rely on pre-trained knowledge rather than actual visual or temporal understanding. (3) Multi-Modal Importance: When provided with full video and subtitle context, however, models show substantial improvements, confirming the critical role of multimodal input in video understanding.
@@
Short TGIF-QA (Jang et al., 2017) 8.5 K 9575 0.05 7.98
MSRVTT-QA (Xu et al., 2017) 72.8 K 2990 0.25 12.45
MV-Bench (Li et al., 2024b) 4.0 K 3641 0.27 16.38
Long Activity-QA (Yu et al., 2019) 8.0 K 800 1.85 24.67
TVQA (Lei et al., 2019) 15.2 K 2179 1.86 67.55
@@
InfiniBench (Ours) 91K 1219 52.59 1068.45
Table 1: Comparison between InfiniBench and existing video understanding benchmarks. InfiniBench has the largest QA pairs and the longest total video durations.
@@ 1 Introduction
Recent progress in Multimodal large language models (MLMMs) has enabled impressive performance on image and short-video understanding by jointly processing visual and textual inputs (Ataallah et al., 2024a; Zhu et al., 2023; Zhang et al., 2023a; Chen et al., 2023; Lin et al., 2023; Liu et al., 2023; Maaz et al., 2023). To push toward richer video comprehension, several models have extended the context length of underlying LLMs, enabling them to process longer videos  (Bai et al., 2023, 2025; Li et al., 2024a). However, current public benchmarks remain limited in both scope and realism. Most are restricted to short clips (a few seconds to minutes), and even larger-scale efforts like LVBench, with one-hour videos, lack narrative complexity and modality diversity (Wang et al., 2024b). In real-world scenarios, long-form videos especially movies and TV shows present fundamentally different challenges. These formats feature rich, structured storytelling with long-range temporal dependencies, evolving character arcs, and intricate causal relationships (Song et al., 2023; Regneri et al., 2013; Rohrbach et al., 2014; Awad et al., 2017, 2018, 2020). Unlike short videos, they demand high-level skills such as intent inference, multi-event linking, and thematic summarization. Humans can easily integrate dispersed cues across modalities and time; current MLMMs cannot.
To bridge this gap, we introduce InfiniBench, a large-scale benchmark purpose-built to evaluate the long-video understanding capabilities of multi-modal models. InfiniBench leverages over 1,000 hours of content drawn from movies and TV series with average video length (52.59 minutes) and the largest set of high-quality, skill-targeted question–answer pairs (91K) among existing benchmarks (see Table 1).
@@
Key Insights from InfiniBench: Despite the scale and capabilities of modern MLMMs, all evaluated models struggle on InfiniBench. For instance, GPT-4o, which is the top-performing commercial model, achieves only 47.1% accuracy on grounding-based tasks, just modestly above the random baseline of 20%. This performance gap highlights how far even the best models are from mastering long-video comprehension.
To dissect model behavior further, we conduct controlled experiments that isolate the role of world knowledge vs. visual input. Surprisingly, models still achieve non-trivial accuracy when given only metadata (e.g., show title and episode number), indicating heavy reliance on pre-trained knowledge. Yet, the most substantial gains up to +15% in accuracy come from providing visual input, confirming that true video understanding requires visual grounding, not just memorized associations. Finally, we observe consistent underperformance across both skill groups, grounding and reasoning-based skills. Reasoning-based tasks remain especially challenging, with even commercial models failing to produce coherent, causal, or context-aware answers. These findings underscore the unique difficulty and diagnostic power of InfiniBench, which surfaces fundamental limitations in today’s multi-modal models and points toward critical future directions.
@@ 2 Related Work
Recently, many multimodal large language models have aimed to extend their context lengths, such as Qwen2.5-VL (Bai et al., 2025), Qwen2-VL (Wang et al., 2024a), and LongVU (Wu et al., 2024), which can process videos over an hour long. This highlights the growing need for effective benchmarks. In contrast, short video benchmarks have been widely explored, covering tasks like moment retrieval (Liang et al., 2024; Lei et al., 2021), action classification (Caba Heilbron et al., 2015; Carreira and Zisserman, 2017; Soomro et al., 2012), video reasoning (Xie et al., 2024; Xiao et al., 2021), and video QA (Xu et al., 2017; Lei et al., 2019; Jang et al., 2017). Among longer video benchmarks, TVQA (Lei et al., 2019) is an early example with over 15K QA pairs across 2K clips (1.86 min avg.), totaling 67.55 hours. More recent datasets like MovieChat-1K(Song et al., 2023) offer 1,000 movie clips (9.4 min avg.) with 14K annotations. CRAFT(Ates et al., 2020) and GazeVQA (Ilaslan et al., 2023) extend to causal and embodied reasoning. Other long-form efforts include MLVU (Zhou et al., 2024), MoVQA (Zhang et al., 2023b), and Video-MME (Fu et al., 2024), though their average durations peak at 17 minutes. LVBench(Wang et al., 2024b) stands out with 1-hour videos, but includes only 103 clips totaling 117 hours. Domain-specific datasets like Ego4D QA on EgoSchema(Mangalam et al., 2023) focus on first-person views but use short (3-minute) clips. Motivated by these gaps, our benchmark evaluates grounding (temporal/visual anchoring) and reasoning (causality, narrative) skills to reflect human-like video understanding. As shown in Table 1, our dataset supports near-hour-long videos and offers the most significant total duration ($\sim$1K hours) and the largest QA-pair count (91K).
@@ 3.3 Data Statistics
InfiniBench ultimately comprises 91K questions over videos averaging 53 minutes, with some running as long as 201 minutes, making it the largest and most challenging benchmark for long-video understanding. InfiniBench includes 923 episodes from six TV shows (Lei et al., 2019) based on the long-form version in GoldFish Ataallah et al. (2024b) and filtered 296 movies that have subtitles from MovieNet (Huang et al., 2020). These movies offer a broad and challenging foundation for benchmarking long-video understanding. Figures 4 and 4 illustrate the per-skill distribution, source breakdown (TVQA (Lei et al., 2019), MovieNet (Huang et al., 2020)), and duration statistics across the dataset. We humanly verified 20% of the data, ensuring fair coverage across TV shows and movies. The remaining 80 % is available for training video models. The verified portion is further divided into test and validation splits, 10% each. All reported results and ablation studies are conducted using the test split.
@@ Appendix C Extended Related Work
Recently, many multimodal large language models have aimed to extend their context lengths, such as Qwen2.5-VL (Bai et al., 2025), Qwen2-VL (Wang et al., 2024a), and LongVU (Wu et al., 2024), which can process videos over an hour long. This highlights the growing need for effective benchmarks. In contrast, short video benchmarks have been widely explored, covering tasks like moment retrieval (Liang et al., 2024; Lei et al., 2021), action classification (Caba Heilbron et al., 2015; Carreira and Zisserman, 2017; Soomro et al., 2012), video reasoning (Xie et al., 2024; Xiao et al., 2021), and video QA (Xu et al., 2017; Lei et al., 2019; Jang et al., 2017). Among longer video benchmarks, TVQA (Lei et al., 2019) is an early example with over 15K QA pairs across 2K clips (1.86 min avg.), totaling 67.55 hours. More recent datasets like MovieChat-1K(Song et al., 2023) offer 1,000 movie clips (9.4 min avg.) with 14K annotations. CRAFT(Ates et al., 2020) and GazeVQA (Ilaslan et al., 2023) extend to causal and embodied reasoning. Other long-form efforts include MLVU (Zhou et al., 2024), MoVQA (Zhang et al., 2023b), and Video-MME (Fu et al., 2024), though their average durations peak at 17 minutes. LVBench(Wang et al., 2024b) stands out with 1-hour videos, but includes only 103 clips totaling 117 hours. Domain-specific datasets like Ego4D QA on EgoSchema(Mangalam et al., 2023) focus on first-person views but use short (3-minute) clips. Motivated by these gaps, our benchmark evaluates grounding (temporal/visual anchoring) and reasoning (causality, narrative) skills to reflect human-like video understanding. As shown in Table 1, our dataset supports near-hour-long videos and offers the most significant total duration ($\sim$1K hours) and the largest QA-pair count (91K).

### v4

@@ Abstract
Understanding long-form videos, such as movies and TV episodes ranging from tens of minutes to two hours, remains a significant challenge for multi-modal models. Existing benchmarks often fail to test the full range of cognitive skills needed to process these temporally rich and narratively complex inputs. Therefore, we introduce InfiniBench, a comprehensive benchmark designed to evaluate the capabilities of models in long video understanding rigorously. InfiniBench offers: (1) Over 1,000 hours of video content, with an average video length of 53 minutes. (2) The largest set of question-answer pairs for long video comprehension, totaling around 87.7K. (3) Eight diverse skills that span both grounding-based (e.g., scene transitions, character actions) and reasoning-based (e.g., deep context understanding, multi-event linking). (4) Rich annotation formats, including both multiple-choice and open-ended questions. We conducted an in-depth evaluation across both commercial (GPT-4o, Gemini 2.0 Flash) and most recent open-source vision-language models such as Qwen2.5-VL, InternVL3.0). Results reveal that: (1) Models struggle across the board: Even the best model, GPT-4o, achieves only 47.1% on grounding-based skills, with most models performing near or just above random chance. (2) Strong reliance on world knowledge: Models achieve surprisingly high scores using only metadata (e.g., video titles), highlighting a tendency to rely on pre-trained knowledge rather than actual visual or temporal understanding. (3) Multi-Modal Importance: When provided with full video and subtitle context, however, models show substantial improvements, confirming the critical role of multimodal input in video understanding.
@@
Short TGIF-QA (Jang et al., 2017) 165.2 K 72.0 K 0.05 60.00
MSRVTT-QA (Xu et al., 2017) 243.6 K 10.0 K 0.25 416.6
MV-Bench (Li et al., 2024b) 4.0 K 3.6 K 0.27 16.38
TVQA (Lei et al., 2019) 152.5 K 21.8K 1.27 461.20
Long Activity-QA (Yu et al., 2019) 58.0 K 5800 3.00 290.00
@@
InfiniBench (Ours) 87.7K 1217 52.59 1066.70
Table 1: Comparison between InfiniBench and existing video QA benchmarks and datasets. InfiniBench contains the largest number of QA pairs for long and very long videos, as well as the longest total video duration.
@@ 1 Introduction
Recent progress in Multimodal large language models (MLMMs) has enabled impressive performance on image and short-video understanding by jointly processing visual and textual inputs (Ataallah et al., 2024a; Zhu et al., 2023; Zhang et al., 2023a; Chen et al., 2023; Lin et al., 2023; Liu et al., 2023; Maaz et al., 2023). To push toward richer video comprehension, several models have extended the context length of underlying LLMs, enabling them to process longer videos  (Bai et al., 2023, 2025; Li et al., 2024a). However, current public benchmarks remain limited in both scope and realism. Most are restricted to short clips (a few seconds to minutes), and even larger-scale efforts like LVBench, with one-hour videos, lack narrative complexity and modality diversity (Wang et al., 2024b). In real-world scenarios, long-form videos especially movies and TV shows present fundamentally different challenges. These formats feature rich, structured storytelling with long-range temporal dependencies, evolving character arcs, and intricate causal relationships (Zhang et al., 2023b; Lei et al., 2019; Song et al., 2023) . Unlike short videos, they demand high-level skills such as intent inference, multi-event linking, and thematic summarization. Humans can easily integrate dispersed cues across modalities and time; current MLMMs cannot.
To bridge this gap, we introduce InfiniBench, a large-scale benchmark purpose-built to evaluate the long-video understanding capabilities of multi-modal models. InfiniBench leverages over 1,000 hours of content drawn from movies and TV series with average video length (52.59 minutes) and the largest set of skill-targeted question–answer pairs (87.7K) among existing long and very long video benchmarks (see Table 1).
@@
Key Insights from InfiniBench: Despite the scale and capabilities of modern MLMMs, all evaluated models struggle on InfiniBench. For instance, GPT-4o, which is the top-performing commercial model, achieves only 47.1% accuracy on grounding-based tasks, just modestly above the random baseline of 20%. This performance gap highlights how far even the best models are from mastering long-video comprehension. To dissect model behavior further, we conduct controlled experiments that isolate the role of world knowledge vs. visual input. Surprisingly, models still achieve non-trivial accuracy when given only metadata (e.g., show title and episode number), indicating heavy reliance on pre-trained knowledge. Yet, the most substantial gains up to +15% in accuracy come from providing visual input, confirming that true video understanding requires visual grounding, not just memorized associations. Finally, we observe consistent underperformance across both skill groups, grounding and reasoning-based skills. Reasoning-based tasks remain especially challenging, with even commercial models failing to produce coherent, causal, or context-aware answers. These findings underscore the unique difficulty and diagnostic power of InfiniBench, which surfaces fundamental limitations in today’s multi-modal models and points toward critical future directions.
@@ 2 Related Work
Recently, many multimodal large language models have aimed to extend their context lengths, such as Qwen2.5-VL (Bai et al., 2025), Qwen2-VL (Wang et al., 2024a), and LongVU (Wu et al., 2024), which can process videos over an hour long. This highlights the growing need for effective benchmarks. In contrast, short video benchmarks have been widely explored, covering tasks like moment retrieval (Liang et al., 2024; Lei et al., 2021), action classification (Caba Heilbron et al., 2015; Carreira and Zisserman, 2017; Soomro et al., 2012), video reasoning (Xie et al., 2024; Xiao et al., 2021), and video QA (Xu et al., 2017; Lei et al., 2019; Jang et al., 2017). Among longer video benchmarks, TVQA (Lei et al., 2019) is an early example with over 152.5 K QA pairs across 21.8 K clips (1.27 min avg.), totaling 461.2 hours. More recent datasets like MovieChat-1K(Song et al., 2023) offer 1,000 movie clips (9.4 min avg.) with 14K annotations. CRAFT(Ates et al., 2020) and GazeVQA (Ilaslan et al., 2023) extend to causal and embodied reasoning. Other long-form efforts include MLVU (Zhou et al., 2024), MoVQA (Zhang et al., 2023b), and Video-MME (Fu et al., 2024), though their average durations peak at 17 minutes. LVBench(Wang et al., 2024b) stands out with 1-hour videos, but includes only 103 clips totaling 117 hours. Domain-specific datasets like Ego4D QA on EgoSchema(Mangalam et al., 2023) focus on first-person views but use short (3-minute) clips. Motivated by these gaps, our benchmark evaluates grounding (temporal/visual anchoring) and reasoning (causality, narrative) skills to reflect human-like video understanding. As shown in Table 1, our dataset supports near-hour-long videos and offers the most significant total duration ($\sim$1K hours) and the largest QA-pair count in terms of long video understanding (87.7K).
@@ 3.3 Data Statistics
InfiniBench ultimately comprises 87.7K questions over videos averaging 53 minutes, with some running as long as 201 minutes, making it the largest and most challenging benchmark for long-video understanding. InfiniBench includes 923 episodes from six TV shows (Lei et al., 2019) based on the long-form version in GoldFish Ataallah et al. (2024b) and filtered 296 movies that have subtitles from MovieNet (Huang et al., 2020). These movies offer a broad and challenging foundation for benchmarking long-video understanding. Figures 4 and 4 illustrate the per-skill distribution, source breakdown (TVQA (Lei et al., 2019), MovieNet (Huang et al., 2020)), and duration statistics across the dataset. We humanly verified 20% of the data, ensuring fair coverage across TV shows and movies. The remaining 80 % is available for training video models. The verified portion is further divided into test and validation splits, 10% each. All reported results and ablation studies are conducted using the test split.
@@ Appendix C Extended Related Work
Recently, many multimodal large language models have aimed to extend their context lengths, such as Qwen2.5-VL (Bai et al., 2025), Qwen2-VL (Wang et al., 2024a), and LongVU (Wu et al., 2024), which can process videos over an hour long. This highlights the growing need for effective benchmarks. In contrast, short video benchmarks have been widely explored, covering tasks like moment retrieval (Liang et al., 2024; Lei et al., 2021), action classification (Caba Heilbron et al., 2015; Carreira and Zisserman, 2017; Soomro et al., 2012), video reasoning (Xie et al., 2024; Xiao et al., 2021), and video QA (Xu et al., 2017; Lei et al., 2019; Jang et al., 2017). Among longer video benchmarks, TVQA (Lei et al., 2019) is an early example with over 152.5 K QA pairs across 21.8 K clips (1.27 min avg.), totaling 461.2 hours. More recent datasets like MovieChat-1K(Song et al., 2023) offer 1,000 movie clips (9.4 min avg.) with 14K annotations. CRAFT(Ates et al., 2020) and GazeVQA (Ilaslan et al., 2023) extend to causal and embodied reasoning. Other long-form efforts include MLVU (Zhou et al., 2024), MoVQA (Zhang et al., 2023b), and Video-MME (Fu et al., 2024), though their average durations peak at 17 minutes. LVBench(Wang et al., 2024b) stands out with 1-hour videos, but includes only 103 clips totaling 117 hours. Domain-specific datasets like Ego4D QA on EgoSchema(Mangalam et al., 2023) focus on first-person views but use short (3-minute) clips. Motivated by these gaps, our benchmark evaluates grounding (temporal/visual anchoring) and reasoning (causality, narrative) skills to reflect human-like video understanding. As shown in Table 1, our dataset supports near-hour-long videos and offers the most significant total duration ($\sim$1K hours) and the largest QA-pair count in terms of long video understanding (87.7K).